OPTICAL-OPTICAL DOUBLE RESONANCE SPECTROSCOPY OF CaOH: INVESTIGATING THE HIGH ENERGY STATES

\underline{M. J. DICK}, Department of Physics, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada; J.-G. WANG, P. M. SHERIDAN and P. F. BERNATH, Department of Chemistry, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada.

The D̃^{2}Σ^{+}(000) state of CaOH has been investigated at high resolution using the technique of optical-optical double resonance spectroscopy. CaOH was synthesized in a Broida-type oven by the reaction of H_{2}O_{2} and calcium vapor. A linear cavity dye laser (∼1 cm^{−1} bandwidth) was used to excite the band heads of the Ã^{2}Π_{1/2} − X̃^{2}Σ^{+} and Ã^{2}Π_{3/2}- X̃^{2}Σ^{+}transitions of CaOH. A Ti:Sapphire ring laser was then used to promote the molecules from the Ã^{2}Π state to the D̃^{2}Σ^{+} state. Assignment and analysis of the high resolution spectra is in progress and a discussion of the rotational and fine structure constants of the D̃^{2}Σ^{+}(000) state of CaOH will be presented. In addition, high resolution studies of other higher-lying states of CaOH using optical-optical double resonance spectroscopy are in progress and preliminary results will be described.